Given that the electrolytic conductivity of a solution containing 3g of calcium chloride in 250cm^-3 aqueous solution is 1.6x10^-3 Sm^-1 at 20°C. Calculate the molar conductivity of the solution at 20°C.
1
Expert's answer
2020-10-27T08:24:22-0400

Number of moles = "\frac{M_{given}}{M_{molecular}}"

"= \frac{3}{111}"

= 0.027 mol

Molarity = (Number of Moles)/Volume

= "\frac{0.027}{0.25}"

= 0.108 molL-1

Molar Conductivity = "\kappa" /(1000 "\times" Molarity in molL-1)

= "\frac{1.6\times10^{-3}}{1000\times0.108}"

= 1.48 "\times" 10-5 Sm2mol-1